OpenStudy (anonymous):

Explain (in your own words) the difference between zero divided by seven and seven divided by zero. Show a way of writing these two problems using mathematical symbols for division. Don’t forget to write the solution to each one!

OpenStudy (anonymous):

there is no difference 7/0=0 0/7=0 the only difference is in fractional for \[\frac{ 7 }{ 0 }\] isnt possible

OpenStudy (anonymous):

\[\frac07=0\div7=0\\\frac70=7\div0\text{ is undefined}\]

OpenStudy (precal):

 that is how I like to remember it........